# Integrability for the spectrum of Jordanian AdS/CFT

### v1

@@ 1 Introduction
Concretely, we start from the differential-operator realisation of the transfer matrix and solve its eigenvalue problem for $J=2$ and arbitrary spin $S$, obtaining closed-form expressions (in $S$) for both the eigenvalues and eigenfunctions as perturbative series in the deformation parameter. We then verify explicitly that the twisted Hamiltonian is diagonal on the resulting $S=0$ and $S=1$ eigenstates and compute their corresponding energies. Building on these results, we propose a Baxter relation for the Jordanian-twisted chain whose functional form is the same as that of the undeformed model (cf. also Driezen:2025dww ). The deformation enters only through several fixed coefficients of the transfer matrix. The $Q$-functions cease to be polynomials, and we propose that regularity in the complex plane replaces polynomiality as the condition selecting physical solutions. Notably, the asymptotics of the $Q$-functions are simpler than in the abelian dipole case Guica:2017mtd . We then show that this framework provides the full $J=2$ spectrum very efficiently (including dependence on the spin $S$) and in perfect agreement with the nontrivial direct diagonalisation results. This thus gives compelling evidence for the validity of the given Baxter framework.
@@ 2.1 Spectral objects
Although the Hamiltonian (3) is invariant under global $\mathfrak{sl}(2,\mathbb{R})$ transformations (since each $h_{j,j+1}$ depends only on the two-site Casimir), its diagonalisation is far from straightforward. To tackle the spectral problem one must instead employ integrability techniques. These introduce an auxiliary space $\mathsf{a}$, typically chosen to be the simplest finite-dimensional representation of $\mathfrak{sl}(2,\mathbb{R})$, i.e. with generators $X_{\mathsf{a}}$ acting through standard matrix operations as
@@
and since the transfer matrix by construction is a polynomial of degree $J$ in $u$, its expansion generates $J$ mutually commuting operators. These operators also commute with the physical Hamiltonian, so that their spectrum provides $J$ conserved quantities. Solving the eigenvalue problem of the transfer matrix therefore simultaneously yields the eigenstates of the Hamiltonian as well as the spectrum of conserved charges encoded in $\hat{\tau}(u)$; for the current model see Faddeev:1994zg ; Derkachov:2002tf .
@@ 2.2 $J=2$ eigenvalue problem
where $\tau^{\star}_{0}$ is determined by the ODE for $g^{\star}(z)$,
@@
Solving the equations order by order in $z$ and in in $\xi M$, we then obtain for the transfer matrix eigenvalue
@@
In practice, we have computed the eigenvalues $\tau^{\star}_{0}$ and eigenfunctions $g^{\star}(z)$ for arbitrary $S$ up to ${\cal O}(\xi^{4}M^{4})$ included, and the Frobenius expansion of each $g_{l}(z)$ up to ${\cal O}(z^{6})$ included. However, above we only present the truncated solutions for brevity. In principle, the expansion can be extended to arbitrary order. For the lowest excitations $S=0,1$, we obtained compact closed-form expressions in $z$, which are presented in appendix A.
@@
The pattern of the expansion of excited states is also noteworthy, with gamma functions of the form $\Gamma(S+\tfrac{2n+1}{2})$ repeatedly appearing, hinting at possible closed-forms for both eigenfunctions and eigenvalues. Furthermore, the analysis clearly shows that each undeformed excited eigenstate admits a smoothly deformed counterpart, with a corresponding deformation of $\tau^{\star}(u)$, demonstrating that the full undeformed spectrum deformes coherently under the Jordanian twist. Finally, it is also instructive to compare the present approach, based on a perturbative expansion in $\xi$, with that of Driezen:2025dww , where regularity in $z$ was imposed first. The two approaches appear to be related by a nontrivial treatment of integration constants: allowing these constants to depend on $\xi$ effectively interchanges the order of expansions and allows to find excited states above the deformed vacuum. Indeed, taking the leading integration constant in eq. (5.22) of Driezen:2025dww to be $g_{0}=\xi^{S}$ for $S=0,1,2$ rather than $g_{0}=1$ prior to demanding regularity at $\xi=0$, reproduces precisely the $S=0,1,2$ deformed eigenstates and eigenvalues presented in this section. However, extending this treatment to arbitrary $S$ remains technically cumbersome, as it would require solving the recurrence relation eq. (5.23) of Driezen:2025dww in closed form such that the choice of integration constants can be tracked. By contrast, expanding directly in $\xi$ as done above provides a more efficient and systematic solution scheme for all $S$.
@@ 3.3 Asymptotics and regularity of $Q$-functions
where the subscript $a=1,2$ indicates that the Baxter equation has two independent solutions with different asymptotics, with exponents $\phi_{1}=\phi,\ \phi_{2}=-\phi$. The true physical $Q$-function will be a particular linear combination of these two asymptotic series. This behaviour immediately shows that the Jordanian $Q$-functions are neither polynomials nor simple exponentials times polynomials (as in the diagonally twisted case), but instead have a much more complicated structure. Note however that these asymptotics are much simpler on the complex curve than in the dipole case.
@@
Our analysis shows that the correct requirement for the Jordanian model is analytic regularity of $Q$, i.e. that the physical solutions are those with no singularities for all finite values of $u$ (except possibly at infinity). This is a crucial part which completes our proposal for how the Baxter equation works in the Jordanian model. In fact, given the special large $u$ asymptotics, it is not even obvious that such solutions exist at all. Remarkably, as we show below, regular solutions do exist, and imposing this condition indeed restricts both $Q$ and the transfer matrix eigenvalue to a discrete set of solutions corresponding to the spin chain eigenstates. This regularity condition coincides with that used for the dipole deformation, where again the asymptotics exclude any kind of simple function for $Q$ (however there the asymptotics are far more complex).
@@ 3.4.1 $J=2$ spectrum
This expression constitutes the full energy spectrum of the $J=2$ Jordanian-twisted spin chain. Notice that each undeformed level labelled by spin $S$ acquires a distinct deformation. However, in the deformed theory, let us stress that $S$ is only a convenient label for the resulting branches of states rather than a conserved quantum number, since for $\xi\neq 0$ the Hamiltonian no longer commuted with the two-site Casimir.
@@ 3.6 The most general 2-parameter deformation
In section 3.3 we compared different deformations that correspond to having different nontrivial coefficients in the transfer matrix. Here we briefly discuss one case not included there – namely, when the leading and subleading coefficients are both fixed to generic values. While it is not yet clear how to actually build the corresponding deformed spin chain, let us explore here what happens to the $Q$-functions.
@@
where $\alpha,\beta$ are arbitrary finite parameters, with $\alpha=\beta=0$ giving the undeformed case. The Jordanian deformation, the dipole deformation, and the diagonal twist deformation [^25] In the diagonal twist case $\beta$ is not actually a free parameter, rather it is fixed by imposing that $Q$ is an exponent times polynomial. are all special cases of this expression (74). Let us show explicitly for $J=2$ how we can solve the Baxter equation as an expansion in $\alpha,\beta$ that we assume to be perturbative small parameters (with their ratio $\alpha/\beta$ kept fixed). We write
@@ 4.4 Semiclassical spectrum and matching
    Let us recall the classical relation (90) found from the spectral curve of the BMN-like string. In Borsato:2022drc it was shown that $\mathbf{Q}$ does not receive any semiclassical corrections, implying that $\mathbf{Q}=2\pi\eta\sqrt{\lambda}m$ also holds semiclassically. Together with the identification (103) we can thus conclude that the string twist charge $\mathbf{Q}$ relates to the spin chain parameter through $\log(1+\xi M)$. This identification is physically well-motivated: Firstly, at the classical level it reduces to $\sqrt{\lambda}m\sim M$, as we found in the Landau-Lifshitz limit, when identifying $\eta=\tfrac{\sqrt{\lambda}}{2\pi}\xi$ (exactly as was needed when mapping the string and spin chain Drinfel’d twist, cf. footnote 36). At the quantum level, one can understand the logarithmic structure as reflecting the non-abelian nature of the Jordanian twist. In particular, this should be contrasted with abelian twists such as the dipole/Schrödinger deformation where the momentum charges map also at the quantum level simply as $\sqrt{\lambda}m\sim M$ Guica:2017mtd ; Ouyang:2017yko .
@@ 5 Conclusions
We proposed a Baxter framework for the Jordanian deformation with intriguing and fully novel structural features compared to previously known cases in the literature (as summarised in section 3.3). In particular, despite the fact that the Yangian (RTT) structure is deformed, we propose that the functional form of the $TQ$-relation and its energy formula remain undeformed. The deformation enters only through fixed coefficients in the transfer matrix, resulting in non-polynomial $Q$-functions with exponential asymptotics. An additional crucial ingredient is that analytic regularity in the complex plane replaces the polynomiality requirement of undeformed chains in order to select spectral $Q$-functions, similar as in Guica:2017mtd .
The direct computation of the $J=2$ spectrum from the Hamiltonian then provides a crucial validation of this framework, as it perfectly matches the spectral results from Baxter across higher spin sectors. Importantly, this agreement supports the expectation that the functional form of the $TQ$-relation is universal for all models based on rational $R$-matrices, as commented on in Guica:2017mtd ; Driezen:2025dww , even when the underlying algebraic structure is substantially modified. We leave for the future how this fits into a more invariant, representation-theoretic point of view on $Q$-operators Chervov:2006xk ; Bazhanov:1996dr ; Bazhanov:2010ts .
@@
In addition, the compatibility between the analyticity conditions for the $Q$-functions of the Jordanian and dipole deformation Guica:2017mtd led us to consider a more general model, combining in spirit both deformations (yet whose physical realisation we leave open). We illustrated for $J=2$ and $S=0$ that its spectrum can also be solved with the Baxter equation.
@@
With the Baxter framework established for the Jordanian chain, we efficiently computed the spectrum to arbitrary length $J$ for the ground and first excited states. This enabled comparison with the string worldsheet theory, conjectured via the work of vanTongeren:2015uha to be the Jordanian HYB-deformation in the $\mathfrak{sl}(2,R)$ sector. In particular, the spectrum of the (nearest-neighbour) twisted chain in the large-$J$ limit matches rather fascinatingly with the semiclassical string spectrum Borsato:2022drc from the algebraic curve to order ${\cal O}(J^{-2})$ included—exactly the same order at which the correspondence holds as in undeformed AdS/CFT before long-range interactions trigger discrepancies Beisert:2005cw . This is particularly striking given that the Jordanian model has very little Noether (super)symmetry. The agreement despite this symmetry reduction suggests that integrability itself appears sufficient to stabilise the correspondence at this order.
@@ Acknowledgements
We thank N. Beisert, S. Ekhammar, G.  Felder, G. Ferrando, N. Gromov, E. Im, D. Serban, C. Thull, B. Vicedo and K. Zarembo for discussions. The work of F.L.-M. was supported by the STFC grant APP69281. The work of S.D. is supported by the Swiss National Science Foundation through the SPF fellowship TMPFP2_224600\_224600 and both A.M. and S.D. acknowledge support by the NCCR SwissMAP.
@@ Appendix A Closed-form $J=2$ eigenstates for $S=0,1$
For the lowest excitations $S=0,1$, and upon appropriate normalisation as explained in the main text, one can directly solve the ODEs in the lowest orders in $\xi$ without having to perform a Frobenius analysis. This thus gives compact analytic expressions for the $J=2$ eigenfunctions in the relative coordinate $z$. For $S=0$, they read

### v2

@@ 1 Introduction
Concretely, we start from the differential-operator realisation of the transfer matrix and solve its eigenvalue problem for $J=2$ and arbitrary spin $S$, obtaining closed-form expressions (in $S$) for both the eigenvalues and eigenfunctions as perturbative series in the deformation parameter. We then verify explicitly that the twisted Hamiltonian is diagonal on the resulting $S=0$ and $S=1$ eigenstates and compute their corresponding energies. Building on these results, we propose a Baxter relation for the Jordanian-twisted chain whose functional form is the same as that of the undeformed model (cf. also Driezen:2025dww ). The deformation enters only through several fixed coefficients of the transfer matrix. The $Q$-functions cease to be polynomials, and we propose that regularity in the complex plane replaces polynomiality as the condition selecting physical solutions. Notably, the asymptotics of the $Q$-functions are simpler than in the abelian dipole case Guica:2017mtd . We then show that this framework provides the full $J=2$ spectrum very efficiently (including dependence on the spin label $S$) and in perfect agreement with the nontrivial direct diagonalisation results. This thus gives compelling evidence for the validity of the given Baxter framework.
@@ 2.1 Spectral objects
While the Hamiltonian (3) is invariant under global $\mathfrak{sl}(2,\mathbb{R})$ transformations (since each $h_{j,j+1}$ depends only on the two-site Casimir), its diagonalisation is far from straightforward. To tackle the spectral problem one can however employ integrability techniques. These introduce an auxiliary space $\mathsf{a}$, typically chosen to be the simplest finite-dimensional representation of $\mathfrak{sl}(2,\mathbb{R})$, i.e. with generators $X_{\mathsf{a}}$ acting through standard matrix operations as
@@
and since the transfer matrix by construction is a polynomial of degree $J$ in $u$, its expansion generates $J$ mutually commuting operators. However, when the auxiliary representation differs from the physical representation, the Hamiltonian itself will not be part of this spectral expansion. Nevertheless, the commuting operators from the transfer matrix also commute with the physical Hamiltonian, and thus their spectrum provides $J$ conserved quantities. Therefore, solving the eigenvalue problem of the transfer matrix will simultaneously yield the eigenstates of the Hamiltonian as well as the spectrum of conserved charges encoded in $\hat{\tau}(u)$; for the current model see Faddeev:1994zg ; Derkachov:2002tf .
@@
Before starting let us already remark that, although we will label the deformed spectrum by the spin number $S$ of the undeformed eigenstates, the deformed eigenstates are not eigenstates of the quadratic Casimir. Rather, we use $S$ only as a book-keeping index labelling deformations of undeformed states.
@@ 2.2 $J=2$ eigenvalue problem
where $\tau^{\star}_{0}$ is determined by the ${\cal O}(u^{0})$ term of $\hat{\tau}^{\star}(u)\Psi^{\star}(z_{1},z_{2})={\tau}^{\star}(u)\Psi^{\star}(z_{1},z_{2})$, which gives an ODE for $g^{\star}(z)$,
@@
Solving the equations order by order in $z$ and in $\xi M$, we then obtain for the transfer matrix eigenvalue
@@
In practice, we have computed the eigenvalues $\tau^{\star}_{0}$ and eigenfunctions $g^{\star}(z)$ for arbitrary $S$ up to ${\cal O}(\xi^{4}M^{4})$ included, and the Frobenius expansion of each $g_{l}(z)$ up to ${\cal O}(z^{6})$ included. However, above we only present the truncated solutions for brevity. In principle, the expansion can be extended to arbitrary order. For the deformations of the lowest excitations $S=0,1$, we obtained compact closed-form expressions in $z$, which are presented in appendix A.
@@
The pattern of the expansion of excited states is also noteworthy, with gamma functions of the form $\Gamma(S+\tfrac{2n+1}{2})$ repeatedly appearing, hinting at possible closed-forms for both eigenfunctions and eigenvalues. Furthermore, the analysis clearly shows that each undeformed excited eigenstate admits a smoothly deformed counterpart, with a corresponding deformation of $\tau^{\star}(u)$, demonstrating that the full undeformed spectrum deforms coherently under the Jordanian twist. Finally, it is also instructive to compare the present approach, based on a perturbative expansion in $\xi$, with that of Driezen:2025dww , where regularity in $z$ was imposed first. The two approaches appear to be related by a nontrivial treatment of integration constants: allowing these constants to depend on $\xi$ effectively interchanges the order of expansions and allows to find excited states above the deformed vacuum. Indeed, taking the leading integration constant in eq. (5.22) of Driezen:2025dww to be $g_{0}=\xi^{S}$ for $S=0,1,2$ rather than $g_{0}=1$ prior to demanding regularity at $\xi=0$, reproduces precisely the $S=0,1,2$ deformed eigenstates and eigenvalues presented in this section. However, extending this treatment to arbitrary $S$ remains technically cumbersome, as it would require solving the recurrence relation eq. (5.23) of Driezen:2025dww in closed form such that the choice of integration constants can be tracked. By contrast, expanding directly in $\xi$ as done above provides a more efficient and systematic solution scheme for all $S$.
@@ 3.3 Asymptotics and regularity of $Q$-functions
where the subscript $a=1,2$ indicates that the Baxter equation has two independent solutions with different asymptotics, with exponents $\phi_{1}=\phi,\ \phi_{2}=-\phi$. The true physical $Q$-function will be a particular linear combination of these two asymptotic series. This behaviour immediately shows that the Jordanian $Q$-functions are neither polynomials nor simple exponentials times polynomials (as in the diagonally twisted case), but instead have a more complicated structure. Note however that on the complex curve these asymptotics are much simpler than in the dipole case.
@@
The analysis of the following subsections will show that the correct requirement for the Jordanian model is analytic regularity of $Q$, i.e. that the physical solutions are those with no singularities for all finite values of $u$ (except possibly at infinity). [^13] This can be seen as the minimal analytic requirement compatible both with polynomiality in the undeformed limit and with the deformed asymptotics, motivated by similar requirements in Guica:2017mtd . This is a crucial part which completes our proposal for how the Baxter equation works in the Jordanian model. In fact, given the special large $u$ asymptotics, it is not even obvious that such solutions exist at all. Remarkably, as we show below, regular solutions do exist, and imposing this condition indeed restricts both $Q$ and the transfer matrix eigenvalue to a discrete set of solutions corresponding to the spin chain eigenstates. This regularity condition coincides with that used for the dipole deformation, where again the asymptotics exclude any kind of simple function for $Q$ (however there the asymptotics are far more complex).
@@ 3.4.1 $J=2$ spectrum
This expression constitutes the full energy spectrum of the $J=2$ Jordanian-twisted spin chain. Notice that each undeformed level labelled by spin $S$ acquires a distinct deformation. However, in the deformed theory, let us stress again that $S$ is only a convenient label for the resulting branches of states rather than a conserved quantum number, since for $\xi\neq 0$ the Hamiltonian no longer commutes with the two-site Casimir.
@@ 3.6 The most general 2-parameter deformation
In section 3.3 we compared different deformations that correspond to having different nontrivial coefficients in the transfer matrix. Here we briefly discuss one case not included there; namely, when the leading and subleading coefficients are both fixed to generic values. While we do not consider the corresponding deformed spin chain here, we will explore what happens to the $Q$-functions under the analytic regularity condition.
@@
where $\alpha,\beta$ are arbitrary finite parameters, with $\alpha=\beta=0$ giving the undeformed case. The Jordanian deformation, the dipole deformation, and the diagonal twist deformation [^26] In the diagonal twist case $\beta$ is not actually a free parameter, rather it is fixed by imposing the additional regularity requirement that $Q$ is an exponent times polynomial. are all special cases of this expression (74). Let us show explicitly for $J=2$ how we can solve the Baxter equation as an expansion in $\alpha,\beta$ that we assume to be perturbative small parameters (with their ratio $\alpha/\beta$ kept fixed). We write
@@ 4.4 Semiclassical spectrum and matching
    Let us recall the classical relation (90) found from the spectral curve of the BMN-like string. In Borsato:2022drc it was shown that $\mathbf{Q}$ does not receive any semiclassical corrections, implying that $\mathbf{Q}=2\pi\eta\sqrt{\lambda}m$ also holds semiclassically. Together with the identification (103) we can thus conclude that the string twist charge $\mathbf{Q}$ relates to the spin chain parameter through $\log(1+\xi M)$. This identification is physically well-motivated: Firstly, at the classical level it reduces to $\sqrt{\lambda}m\sim M$, as we found in the Landau-Lifshitz limit, when identifying $\eta=\tfrac{\sqrt{\lambda}}{2\pi}\xi$ (exactly as was needed when mapping the string and spin chain Drinfel’d twist, cf. footnote 37). At the quantum level, one can understand the logarithmic structure as reflecting the non-abelian nature of the Jordanian twist. In particular, this should be contrasted with abelian twists such as the dipole/Schrödinger deformation where the momentum charges map also at the quantum level simply as $\sqrt{\lambda}m\sim M$ Guica:2017mtd ; Ouyang:2017yko .
@@ 5 Conclusions
We proposed a Baxter framework for the Jordanian deformation with intriguing and fully novel structural features compared to previously known cases in the literature (as summarised in section 3.3). In particular, despite the fact that the Yangian (RTT) structure is deformed, we propose that the functional form of the $TQ$-relation and its energy formula remain undeformed. The deformation enters only through fixed coefficients in the transfer matrix, resulting in non-polynomial $Q$-functions with exponential asymptotics. An additional crucial ingredient is that analytic regularity in the complex plane replaces the polynomiality requirement of undeformed chains in order to select spectral $Q$-functions, similar as in Guica:2017mtd . This proposal is supported by several highly non-trivial tests. The direct computation of the $J=2$ spectrum from the Hamiltonian provides a first crucial validation, as it perfectly matches the spectral results from Baxter across higher spin sectors. We also show more generally that imposing regularity of the $Q$-function fixes the transfer matrix eigenvalue (and the $Q$-function itself). In fact it is non-trivial that a regular solution for the $Q$-function exists at all, further providing strong evidence in favour of our proposal. Importantly, this also supports the expectation that the functional form of the $TQ$-relation is universal for all models based on rational $R$-matrices, as commented on in Guica:2017mtd ; Driezen:2025dww , even when the underlying algebraic structure is substantially modified. [^39] Another potentially useful argument is the observation that Drinfel’d twists reduce to the identity in one-dimensional representations. This fact is relevant because the form of the $TQ$ relation (its $(u\pm i/2)^{J}$ coefficients) is expected to be fixed by transfer matrices with one-dimensional auxiliary spaces (see Bazhanov:2010ts or Frenkel:2013uda ), in which the Drinfel’d twist would never appear, even for a Drinfel’d twisted $R$-matrix/RTTRTT algebra. In the future it would be highly interesting to rigorously derive the Baxter equation and the regularity condition from a full operatorial construction of $Q$-functions, e.g. along the lines of Chervov:2006xk ; Bazhanov:1996dr ; Bazhanov:2010ts or Derkachov:2001yn ; Derkachov:2002tf .
@@
The compatibility between the analyticity conditions for the $Q$-functions of the Jordanian and dipole deformation Guica:2017mtd led us further to consider a more general model, combining in spirit both deformations (yet whose physical realisation we leave open). We illustrated for $J=2$ and $S=0$ that its spectrum can also be solved with the Baxter equation.
@@
With the Baxter framework provided for the Jordanian chain, we efficiently computed the spectrum to arbitrary length $J$ for the ground and first excited states. This enabled comparison with the string worldsheet theory, conjectured via the work of vanTongeren:2015uha to be the Jordanian HYB-deformation in the $\mathfrak{sl}(2,R)$ sector. In particular, the spectrum of the (nearest-neighbour) twisted chain in the large-$J$ limit matches rather fascinatingly with the semiclassical string spectrum Borsato:2022drc from the algebraic curve to order ${\cal O}(J^{-2})$ included—exactly the same order at which the correspondence holds as in undeformed AdS/CFT before long-range interactions trigger discrepancies Beisert:2005cw . This is particularly striking given that the Jordanian model has very little Noether (super)symmetry. The agreement despite this symmetry reduction suggests that integrability itself appears sufficient to stabilise the correspondence at this order.
@@ Acknowledgements
We thank N. Beisert, S. Ekhammar, G. Felder, G. Ferrando, N. Gromov, E. Im, D. Serban, C. Thull, A. Tseytlin, B. Vicedo and K. Zarembo for discussions, and K. Zarembo for comments on the draft. The work of F.L.-M. is supported by the STFC grant APP69281. The work of S.D. is supported by the Swiss National Science Foundation through the SPF fellowship TMPFP2_224600\_224600 and both A.M. and S.D. acknowledge support by the NCCR SwissMAP.
@@ Appendix A Closed-form $J=2$ eigenstates for $S=0,1$
For deformations of the lowest excitations $S=0,1$, and upon appropriate normalisation as explained in the main text, one can directly solve the ODEs in the lowest orders in $\xi$ without having to perform a Frobenius analysis. This thus gives compact analytic expressions for the $J=2$ eigenfunctions in the relative coordinate $z$. For $S=0$, they read